Kirk Minerals processes materials extracted from mines. The most common raw material that it processes results in three joint products: Spock, Uhura, and Sulu. Each of these products can be sold as is, or each can be processed further and sold for a higher price. The company incurs joint costs of $178,600 to process one batch of the raw material that produces the three joint products. The following cost and sales information is available for one batch of each product.

Sales Value at
Split-Off Point
Allocated
Joint Costs
Cost to Process
Further
Sales Value of
Processed Product
Spock $209,500 $39,200 $110,100 $300,900
Uhura 299,900 59,100 85,100 400,900
Sulu 454,000 80,300 249,500 800,700


Determine the incremental profit or loss that each of the three joint products. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Spock Uhura Sulu
Incremental profit (loss) $ $ $



Indicate whether each of the three joint products should be sold as is, or processed further.

Spock process furthersold as is
Uhura process furthersold as is
Sulu process furthersold as is
Answer #1

Answer-

Spock Uhura Sulu
Incremental Profit (Loss) ($18700) $15900 $97200

Explanation-

Kirk Minerals
Statement Of Incremental Profit/(loss)
Product Sale value at Split-Off-Point Sale value of processed product Incremental sale Cost to further process Incremental profit/(loss)
$ $ $ $ $
(a) (b) (c=b-a) (d) (e=c-d)
Spock 209500 300900 91400 110100 -18700
Uhura 299900 400900 101000 85100 15900
Sulu 454000 800700 346700 249500 97200

Answer- Spock = Sold as is

Uhuru= Process further

Sulu = Process further